OpenStudy (anonymous):

Periodic Functuions. A periodic Function f(x) has a period of 12. If f(7)=-2 and f(11)=9, determine the value of f(43). Please explain thoroughly.

hartnn (hartnn):

since the function is periodic with period 12, f(43) = f(43-12) = f(31), did u get this ?

OpenStudy (anonymous):

I think so. So if for example in the same context if the question was f(79) then it would be f(79 - 12) = f(67) ?

hartnn (hartnn):

yes.

hartnn (hartnn):

f(31) = f(31-12) = f(31-12-12) and so on...

OpenStudy (anonymous):

Woah. Why would you subtract 12 twice or more times? And what do we figure out when we do that?

hartnn (hartnn):

because for every interval of 12, the function has same value.... so if i add or subtract 12, the value of function is same \(f(x\pm 12 )= f(x)\)

hartnn (hartnn):

more specifically \(f(x\pm n*12 )= f(x)\)

hartnn (hartnn):

n can be any natural number .

hartnn (hartnn):

so, f(43) = f(43-12) = f(31) = f(31-12) =.... do this till u get something you know the value of.

OpenStudy (anonymous):

not quite sure im understanding.

hartnn (hartnn):

f(43) = f(31) u got this ?

hartnn (hartnn):

no?

hartnn (hartnn):

well, i'll write out steps, see if u get it \(f(43)=f(43-12)=f(31)=f(31-12)=f(19)=f(19-12)=f(7)=-2\)

OpenStudy (anonymous):

Okay I got everything except the f(7)

OpenStudy (anonymous):

Because 19-12 equals 7 but, then 7-12 is equal to -5

hartnn (hartnn):

but u already have f(7) = -2 so u get f(43) = -2

hartnn (hartnn):

and yes, f(-5) also equals -2

OpenStudy (anonymous):

sotheanswer coulve been -5 or -2?

hartnn (hartnn):

no, not -5 . f(43) = f(7) = -2

OpenStudy (anonymous):

hmm. Can we try another example please? say f(-1)?

hartnn (hartnn):

for the same periodic function ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

ok, f(-1) = f(-1+12) = ?

OpenStudy (anonymous):

11

OpenStudy (anonymous):

So it can be plus or minus?

hartnn (hartnn):

its f(11) yes, + or -12 ...

hartnn (hartnn):

f(-1) = f(11) = 9

OpenStudy (anonymous):

Okay. I have an idea of what youre talking about. I'll try this one and can you see if it is correct for me? Same function.

OpenStudy (anonymous):

f(95)

hartnn (hartnn):

yes, try it ..

OpenStudy (anonymous):

Would it be f(95)=f(11)=9 so far am i correct?

hartnn (hartnn):

yes, ver correct :)

OpenStudy (anonymous):

Okay! Thank you once again so much for your patience and time!! Im going to best response all your answers!

OpenStudy (anonymous):

Thats impossible. Lol. Sorry but thankyou!

hartnn (hartnn):

lol. welcome ^_^ and yes, thats not required and not possible...